Bent Gregorian Technical Observing with LUCIFER:
UT 2009 October 29

Observer: DThompson (@LBT)
Telescope Operator: ACeranski
SW support: CBiddick (on call)

Summary

  • Mirror 12C warmer than outside at beginning of the night
  • No observing

Preparations

Details

LUCIFER is not yet copying data to archive, but /newdata mounted to /archive on the lucifer computer. Requires a code change, they will do so at their next recompile.

Calibration unit found to be in the beam. Walter and Marcus moved it out. Calibration unit is unavailable for me to use during this run.

UT23:45 Finished startup with Marcus' help.

23:50 ~90% cloudy. Tout=-10C, Tmirror=0.7C. Wind gusting to 20.5m/s. LBT weather station temperature sensor is flatlined (broken or disconnected?).

01:10 Send preset from LUCIFER for BS9183 ACTIVE PA=254. Failed because M1 was in a panicked state.
01:23 Resent, failed as expected when no guide star was found, but everything else worked.

01:34 Aaron is opening a back vent while the enclosure faces the wind to start cooling the mirror. Drives will be off during this time. Wind gusts to 15m/s. Telescope at horizon.

01:39 Set up to take 1 5s dark manually, looks OK. Images are in /data/20091029/luci.20091029.NNNN.fits (using local date for consistency with commissioning usage, should switch to UT).

02:34 Take two 5x5s darks saved as cubes. Did not click "set new value" so 0002 is just 1x5s. Files 0003 and 0004 are 40MB (2048x2048x5).

02:55 Take two 2x300s darks saved as cubes (0005 and 0006).

03:25 Aaron reports it is clear outside, but M1 is still >10C warm

03:31 Execute darks.txt script. It takes (all o2dcr):
  • 2 "flat"s using setup in script (mistake) 0007, 0008
  • 2 1x2.5s darks 0009, 0010
  • 2 1x5.0s darks 0011, 0012
  • 2 1x25.0s darks 0013, 0014
  • 2 1x125.0s darks 0015, 0016
  • 2 1x600s darks 0017, 0018

I am getting GEIRS Server errors stating "Error saving current buffer to file!" during the above script, but files are being saved to disk and they look OK. Also got an "Unable to get instrument Pressures" error from the Supervisor. The LUCIFER Status window remains green.

04:03 Mirror -1.9C, outside -13.5C

04:06 Switch to K+clear filters. Image 0019 shows AGw probe in the beam. A "home -m 3 -u 3" exited with an error, but probe did move. Will restart AGw and re-home.

05:05 3s K-band exposure inside dome gives ~3000 counts in the center of LUCIFER image, ~10K in 10s. Script Kdomes.txt executed to look at ambient light. Forgot to change the date and rootname from the test script, so they are going into /data/luci/20091027 (I cannot move them back, the observer account does not have privileges). All are single exposure o2dcr images:
  • 2 1x3.0s flats (ambient thermal) 0008, 0009
  • 2 1x6.0s flats (ambient thermal) 0010, 0011
  • 2 1x9.0s flats (ambient thermal) 0012, 0013
  • 2 1x15.0s flats (ambient thermal) 0014, 0015
  • 2 1x25.0s flats (ambient thermal) 0016, 0017
  • 2 1x40.0s flats (ambient thermal) 0018, 0019
  • 2 1x60.0s flats (ambient thermal) saturating in center of image 0020, 0021

Changed destination, rootname, exposure time in script and re-executed.
  • 2 1x2.0s flats (ambient thermal) 0025, 0026
  • 2 1x3.0s flats (ambient thermal) 0027, 0028
  • 2 1x5.0s flats (ambient thermal) 0029, 0030
  • 2 1x8.0s flats (ambient thermal) 0031, 0032
  • 2 1x12.0s flats (ambient thermal) 0033, 0034
  • 2 1x20.0s flats (ambient thermal) 0035, 0036
  • 2 1x35.0s flats (ambient thermal) 0037, 0038
  • 2 1x50.0s flats (ambient thermal) 0039, 0040

Note for Marcus/Walter: headers will need to have a keyword to ID the partner (and probably the PI) for access control in the so that data are appropriately handled. We can discuss this with Riccardo.

Note: editing scripts by hand at 10,000 feet not a good idea. The script syntax is simple enough to understand, but this opens you up to mistakes and thus inefficiencies.

06:00 rising humidity is forcing us to close and stop mirror ventilation. Tout=-13.5, Tmirror=-3.5.

08:48 back to ventilating some time ago. Down to -4.6C on SXM1, -13C outside.

09:50 Stopping mirror ventilation (M1 and the duct temps were the same) at -4.9C. Humidity still high so we cannot reopen to continue cooling.

11:36 Still too humid to open. M1 at -4.9, outside -12.8

Observing Plan

Closed dome
  • Test writing LUCIFER data to archive1 /newdata disk
  • Verify ingestion into archive, propagation to Tucson and on
  • Familiarization with LUCIFER UI

* Correct initial collimation/pointing with IE/CA

* Run dice5 pattern(s) to check that things still work as in May/June

* Check open-loop tracking performance (rms on off-axis GS vs. rms on LUCIFER) as reference data for comparison to post-Nov-HBS work and this has not been checked in a while.

* Collect astrometric field data to check telescope scale at colder temps

* Check guide star FWHM/rms vs LUCIFER image quality (closed loop DIQ)

* Make new collimation model, implement (optional if it looks good)
  • Derive and implement before pointing model obs

* Make new pointing model, implement (if coll model good)
  • Have TO start the pointing log!
  • Check first position in AZ > 360 degrees wrap
  • Use ARC method to measure AGw/Rotator center
  • 28 points are sufficient, more is always better
  • Can use NB filter in LUCIFER, the pointing stars are ~7.5 mag
  • needs stable temps!
  • May need to do several short PMs as temps are changing rapidly

* Take Field aberration measurements if seeing is good (time-series, stability measurements)

* Demonstrate guiding to 18th mag (Tel=GUIDE)
* WFS on 18th mag. Will likely need to set the wfs max exposure time to 120s
* Quantify the guide/wfs limits with mag and seeing vs. LUCIFER DIQ

-- DavidThompson - 26 Oct 2009
Topic revision: r8 - 29 Oct 2009, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback